OpenStudy (anonymous):

solve the trigonomic equation where 0 is less than or equal to x which is less than 2pi : sin^2 2x+3sin2x+2=0

myininaya (myininaya):

let u=sin(2x)

myininaya (myininaya):

u^2+3u+2=0

OpenStudy (anonymous):

(sin2x+1)(sin2x+2)=0 sin2x=-1 only possible s0 x=--pi

myininaya (myininaya):

(u+1)(u+2)=0 u=-1 u=-2

myininaya (myininaya):

but u=sin(2x) so sin(2x)=-1 and sin(2x)=-2 and sin stays between -1 and 1 so yes the second one isn't possible good job dipan

OpenStudy (anonymous):

thanks

myininaya (myininaya):

remember sin(3pi/2)=-1 so we need 2x=3pi/2 just solve for x

myininaya (myininaya):

x=3pi/4